OpenStudy (anonymous):

I am sorry to ask but i really need help. Combine like terms. 14z – 27z

jimthompson5910 (jim_thompson5910):

The variable terms are both the same (both have just a single z in them), so you just subtract the coefficients: 14 - 27 = -13 Then you tack on the variable z to get the final answer of -13z

OpenStudy (hamsandwich):

No need to be sorry :p just subtract 27 from 14 and then put a z at the end. 14 - 27 = z, so it's -13z :D

OpenStudy (anonymous):

Thank you guys alot

OpenStudy (anonymous):

Can you help me with this one is it the same as above Combine like terms. 8 + 16r – 9r

jimthompson5910 (jim_thompson5910):

which 2 terms are the like terms?

OpenStudy (anonymous):

16 and 9

jimthompson5910 (jim_thompson5910):

close, but the terms are 16r and -9r

jimthompson5910 (jim_thompson5910):

the coefficients are 16 and -9 so subtract to get: 16 - 9 = 7 that means 16r - 9r = 7r

jimthompson5910 (jim_thompson5910):

meaning 8 + 16r – 9r simplifies to 8 + 7r

OpenStudy (anonymous):

so then 7r+6 and that is 13

jimthompson5910 (jim_thompson5910):

no you cannot combine 7r and 6 to get 13

OpenStudy (anonymous):

oh it would be 7r+8

jimthompson5910 (jim_thompson5910):

7r and 6 are NOT like terms

OpenStudy (anonymous):

sry i mis typed 6 not 8

jimthompson5910 (jim_thompson5910):

8 + 7r is the same as 7r + 8, so yes

OpenStudy (anonymous):

thank you again

jimthompson5910 (jim_thompson5910):

yw